# PAC-Chernoff Bounds: Understanding Generalization in the Interpolation Regime

### v3

@@ B.4 Invariances (go back to Section 6)
See 7.1
Proof  Due to fact that $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$ and to Theorem 4.1, we have that with high probability $1-\delta$ over $D\sim\nu^{n}(y,x)$,
@@
See 7.2
Proof  Under the isoperimetry assumption, we have ${\cal I}_{\boldsymbol{\theta}}(a)\leq\frac{da^{2}}{2cLip({\boldsymbol{\theta}})^{2}}$. Due to fact that $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$ and to Theorem 4.1, we have that with high probability $1-\delta$ over $D\sim\nu^{n}(y,x)$,
@@
See 7.3
Proof  If the loss is Lipschitz continuous with constant $M$, $\forall y,x,{\boldsymbol{\theta}}\quad\|\nabla_{{\boldsymbol{\theta}}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\|^{2}_{2}\leq M$. Then, $J_{{\boldsymbol{\theta}}}(\lambda)$ verifies
@@
See 7.5
Proof  We can apply Theorem 4.1 on ${\boldsymbol{\theta}}$ assuming that ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}^{\prime}$, because we have that ${\boldsymbol{\Theta}}\subset{\boldsymbol{\Theta}}^{\prime}$. In consequence, with h.p., we have
@@
Then, we can negate the inverse statement using $\neg$ notation as $\neg\Big{[}{\cal I}_{\boldsymbol{\theta}}(a)\geq{\cal I}_{{\boldsymbol{\theta}}^{\prime}}(a)\Big{]}$. Then we have that ${\boldsymbol{\theta}}$ is not $\beta$-smoother than ${\boldsymbol{\theta}}^{\prime}$ for any $\beta\geq a$. Which is equivalent to say that, ${\boldsymbol{\theta}}$ is not $\beta$-smoother than ${\boldsymbol{\theta}}^{\prime}$ for any $\beta$ such that ${\cal I}_{\boldsymbol{\theta}}(\beta)\geq\frac{1}{n}\ln\frac{k^{p^{\prime}}}{\delta}$.
See 7.4
Proof  If ${\boldsymbol{\theta}}^{\prime}$ is $\beta$-smoother than ${\boldsymbol{\theta}}$, by Definition 3, $\forall a\in(0,\beta]\quad{\cal I}_{{\boldsymbol{\theta}}^{\prime}}(a)\geq{\cal I}_{\boldsymbol{\theta}}(a)$ where $\beta={\cal I}^{-1}_{{\boldsymbol{\theta}}^{\prime}}\left(\frac{1}{n}\ln\frac{k^{p^{\prime}}}{\delta}\right)$. Then, we have that ${\cal I}^{-1}_{{\boldsymbol{\theta}}^{\prime}}\left(s\right)\leq{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)$ for $s={\cal I}_{{\boldsymbol{\theta}}^{\prime}}(\beta)=\mathcal{I}_{{\boldsymbol{\theta}}^{\prime}}\Big{(}{\cal I}^{-1}_{{\boldsymbol{\theta}}^{\prime}}\left(\tfrac{1}{n}\ln\tfrac{k^{p^{\prime}}}{\delta}\right)\Big{)}=\tfrac{1}{n}\ln\tfrac{k^{p^{\prime}}}{\delta}$. Thus,

### v4

@@ B.5 Over-Parameterization (go back to Section 8)
See 8.1
Proof.
Due to fact that $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$ and to Theorem 4.1, we have that with high probability $1-\delta$ over $D\sim\nu^{n}(y,x)$,
@@
See 8.2
Proof.
Under the isoperimetry assumption, we have ${\cal I}_{\boldsymbol{\theta}}(a)\leq\frac{da^{2}}{2cLip({\boldsymbol{\theta}})^{2}}$. Due to fact that $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$ and to Theorem 4.1, we have that with high probability $1-\delta$ over $D\sim\nu^{n}(y,x)$,
@@
See 8.3
Proof.
If the loss is Lipschitz continuous with constant $M$, $\forall y,x,{\boldsymbol{\theta}}\quad\|\nabla_{{\boldsymbol{\theta}}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\|^{2}_{2}\leq M$. Then, $J_{{\boldsymbol{\theta}}}(\lambda)$ verifies
@@
See 5.2
Proof.
We can apply Theorem 4.1 on ${\boldsymbol{\theta}}$ assuming that ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}^{\prime}$, because we have that ${\boldsymbol{\Theta}}\subset{\boldsymbol{\Theta}}^{\prime}$. In consequence, with h.p., we have
@@
Then, we can negate the inverse statement using $\neg$ notation as $\neg\Big{[}{\cal I}_{\boldsymbol{\theta}}(a)\geq{\cal I}_{{\boldsymbol{\theta}}^{\prime}}(a)\Big{]}$. Then we have that ${\boldsymbol{\theta}}$ is not $\beta$-smoother than ${\boldsymbol{\theta}}^{\prime}$ for any $\beta\geq a$. Which is equivalent to say that, ${\boldsymbol{\theta}}$ is not $\beta$-smoother than ${\boldsymbol{\theta}}^{\prime}$ for any $\beta$ such that ${\cal I}_{\boldsymbol{\theta}}(\beta)\geq\frac{1}{n}\ln\frac{k^{p^{\prime}}}{\delta}$. ∎
See 5.1
Proof.
If ${\boldsymbol{\theta}}^{\prime}$ is $\beta$-smoother than ${\boldsymbol{\theta}}$, by Definition 3, $\forall a\in(0,\beta]\quad{\cal I}_{{\boldsymbol{\theta}}^{\prime}}(a)\geq{\cal I}_{\boldsymbol{\theta}}(a)$ where $\beta={\cal I}^{-1}_{{\boldsymbol{\theta}}^{\prime}}\left(\frac{1}{n}\ln\frac{k^{p^{\prime}}}{\delta}\right)$. Then, we have that ${\cal I}^{-1}_{{\boldsymbol{\theta}}^{\prime}}\left(s\right)\leq{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)$ for $s={\cal I}_{{\boldsymbol{\theta}}^{\prime}}(\beta)=\mathcal{I}_{{\boldsymbol{\theta}}^{\prime}}\Big{(}{\cal I}^{-1}_{{\boldsymbol{\theta}}^{\prime}}\left(\tfrac{1}{n}\ln\tfrac{k^{p^{\prime}}}{\delta}\right)\Big{)}=\tfrac{1}{n}\ln\tfrac{k^{p^{\prime}}}{\delta}$. Thus,